# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v3

@@ Abstract
Text-guided diffusion models make a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts. Recent works introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pretrained diffusion models for audio modification. Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy. Moreover, existing editing methods fail to achieve effective complex non-rigid music editing while maintaining essential content preservation and high editing fidelity. To counteract these issues, we introduce the Disentangled Inversion technique to disentangle the diffusion process into triple branches, rectifying the deviated path of the source branch caused by DDIM inversion. In addition, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention control and cross-attention control with an intermediate Harmonic Branch to progressively achieve the desired harmonic and melodic information in the target music. Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing while safeguarding content integrity. To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across ten distinct editing categories. This facilitates both zero-shot and instruction-based music editing tasks. Our method achieves unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques. [^1] The project webpage is available at https://MEDIC-Edit.github.io/.
1 Introduction
@@
Text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have made great progress in audio generation (Evans et al., 2024a, b), leveraging their impressive capability for realistic and varied outputs. In particular, these models (Liu et al., 2023c; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes for specific textual prompts. Early audio editing strategies rely on training models from scratch (Copet et al., 2023; Agostinelli et al., 2023) or test-time optimization (Paissan et al., 2023; Plitsis et al., 2024), hampered by intensive computational demands. Recent works (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) and Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) inversion techniques, but challenges remain.
@@
Key among these challenges are maintaining fidelity of editing - ensuring the editing aligns with the provided instructions - and essential content preservation, ensuring that the unaltered musical attributes in the target prompts remain unchanged. However, balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving sub-optimal for conditional diffusion models (Mokady et al., 2023). Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space. However, this comes at the expense of reduced modification capabilities due to noise reduction.
@@
In this work, we rigorously examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis indicates that while techniques like DDIM inversion provide an editable foundation for audio synthesis, they lack precision and may compromise the integrity of the original audio. The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently not met during text-conditional editing. Consequently, this leads to distortions during the inversion. And the implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) aims to improve text adherence. However, it inadvertently amplifies the accumulated errors from the inversion process. Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has shown promise in achieving high fidelity and essential content preservation. For instance, MusicMagus (Zhang et al., 2024) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks. Nevertheless, these methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 1. We introduce an innovative Disentangled Inversion Control technique to bridge this gap. This technique has two main components: Harmonic Attention Control and Disentangled Inversion. Cross-attention control (Hertz et al., 2022) and mutual self-attention control (Cao et al., 2023) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively. However, blindly combining these two approaches sequentially for music editing can result in sub-optimal outcomes, particularly in the original dual-branch setup, where it struggles with global attention refinement. To address this challenge, we introduce an intermediate branch called the Harmonic Branch, designed to modify both rigid and non-rigid attributes in music progressively. Furthermore, we disentangle the diffusion process into triple branches, correcting the deviation path caused by CFG in the source branch, which affects the essential content preservation. The other branches remain unchanged to ensure the highest possible edit fidelity.
@@
Due to a lack of standardized benchmarks in audio editing, we introduce ZoME-Bench. The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories across rigid and non-rigid tasks. Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing. Experimental results on ZoME-Bench indicate that MEDIC outperforms baselines, achieving significant improvements in essential content preservation and editing fidelity. Additionally, MEDIC demonstrates state-of-the-art performance in the variable-length music editing settings of the MusicDelta dataset.
@@
Our contributions can be summarized as follows. 1) We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks. 2) An Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing. 3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information. 4) We build a new benchmark for music editing, named ZoME-Bench, which supports both zero-shot and instruction-based music editing.
2 Related Works
@@
Text-based Audio editing. The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt. Existing methodologies for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction. Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load. The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020; Zhang et al., 2024) inversion strategies to extract diffusion noise vectors that match the source signal. Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework. In this work, we propose a new inversion technique named Disentangled Inversion Control. This technique aims to achieve accurate editing while preserving structural information.
@@
Inversion Techniques. The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021). While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023), its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges. For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during editing. Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024). Optimization-based inversion methods using specific latent variables have recently gained popularity (Ju et al., 2023; Kawar et al., 2023). These are designed to minimize accumulated errors stemming from the DDIM inversion. Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming. Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
@@
Figure 1: Comparisons of our method with two-branch inversion techniques, including DDIM Inversion and MusicMagus. (a) Framework of DDIM Inversion, showing configurations with and without classifier-free guidance. (b) Framework of MusicMagus, which incorporates cross-attention control. (c) Framework of our method, featuring disentangled inversion control. (d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch approach.
3 Disentangled Inversion Control
3.1 Problem Definition and BenchMark Construction
@@
Despite significant work in text-to-audio generation models, particularly with the emergence of latent diffusion models (LDM), research on zero-shot music editing remains limited. Zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to synthesize the desired music, denoted as ${\bm{x}}^{tgt}_{0}$. This synthesized music should align with the target edited text prompt ${\mathcal{P}}^{*}$, which is directly modified from the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt ${\mathcal{P}}$. We compress source audio signals into latent ${\bm{z}}^{src}_{0}$ for inversion.
To systematically validate our proposed method as a plug-and-play strategy for editing models, compare our method with existing zero-shot music editing methods, and compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset named ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps (Doh et al., 2023), spanning ten different editing types that include both rigid and non-rigid modifications. Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
Additionally, we include annotations relevant to attention control, such as blended words, to facilitate detailed evaluations. Further details about our benchmark can be found in Appendix B.
3.2 Motivation

### v4

@@ Abstract.
Text-guided diffusion models revolutionize audio generation by adapting source audio to specific text prompts. However, existing zero-shot audio editing methods such as DDIM inversion accumulate errors across diffusion steps, reducing the effectiveness. Moreover, existing editing methods struggle with conducting complex non-rigid music edits while maintaining content integrity and high fidelity. To address these challenges, we propose MEDIC, a novel zero-shot music editing system based on innovative Disentangled Inversion Control (DIC) technique, which comprises Harmonized Attention Control and Disentangled Inversion. Disentangled Inversion disentangles the diffusion process into triple branches to rectify the deviated path of the source branch caused by DDIM inversion. Harmonized Attention Control unifies the mutual self-attention control and the cross-attention control with an intermediate Harmonic Branch to progressively generate the desired harmonic and melodic information in the target music. We also introduce ZoME-Bench, a comprehensive music editing benchmark with 1,100 samples covering ten distinct editing categories. ZoME-Bench facilitates both zero-shot and instruction-based music editing tasks. Our method outperforms state-of-the-art inversion techniques in editing fidelity and content preservation. The code and benchmark will be released. Audio samples are available at https://melody-edit.github.io/.
Zero-shot Music Editing, Inversion Techniques, Diffusion Models
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755377isbn: 979-8-4007-2035-2/2025/10ccs: Applied computing Sound and music computingccs: Computing methodologies Natural language generation
{subcaptionblock}
0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
Figure 1. Spider chart for CLAP score comparisons.
Figure 2. Spider chart for FAD score comparisons.
Figure 3. Comprehensive zero-shot music editing performance evaluation on the ZoME-bench. We present spider charts of CLAP scores (higher is better) corresponding to editing fidelity and FAD scores (lower is better) corresponding to source-anchored attribute consistency across 10 editing tasks (e.g., change genre) for DDPM-Friendly (manor2024zero), DDIM Inversion (song2020denoising), MusicMagus (zhang2024musicmagus), and our MEDIC with Disentangled Inversion Control.
1. Introduction
@@
Text-guided diffusion models (song2020denoising; song2020improved; peebles2023scalable) have made great progress in audio generation (evans2024fast; evans2024long), leveraging their impressive capability of generating realistic and varied outputs. These models (liu2023audioldm2; huang2023makeanaudio; liu2024audiolcm) provide the foundation for prompt-based music editing, offering new opportunities to modify audio landscapes for specific text prompts. Early music editing strategies rely on training models from scratch (copet2023simple; agostinelli2023musiclm) or test-time optimization (paissan2023audio; plitsis2024investigating), hence they are hampered by intensive computational demands. Recent works (manor2024zero; zhang2024musicmagus) have advanced zero-shot music editing through Denoising Diffusion Implicit Models (DDIM) (song2020denoising) and Denoising Diffusion Probabilistic Models (DDPM) (ho2020denoising) inversion techniques, but challenges remain.
@@
There are two objectives for music editing, that is, edit fidelity - ensuring the editing aligns with the provided instructions - and essential content preservation - maintaining specified musical properties from the source prompt while modifying only designated attributes. For instance, when transforming a melancholic piano composition into an electric guitar arrangement, edit fidelity demands accurate instrument substitution while essential content preservation ensures sustained emotional resonance through retained tempo and recurring minor chord sequences. Balancing these two objectives poses a great challenge since it involves a careful exchange of information between the source and target branch in diffusion processes; however, existing inversion methods such as DDIM prove sub-optimal for conditional diffusion models to address this challenge (mokady2023null). Enhanced versions of edit-friendly DDPM inversion (hubermanspiegelglas2024edit) make strides in content preservation by imprinting the source onto the noise space. However, this method comes at the expense of reduced modification capabilities due to noise reduction.
In this work, we first examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis indicates that while techniques such as DDIM inversion provide a foundation for audio editing, they lack precision and may compromise the integrity of the original audio. The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently violated during text-conditional editing. This issue leads to distortions during the inversion. Although the implementation of Classifier-free Guidance (CFG) (classifier-free) aims to improve text adherence, CFG inadvertently amplifies the accumulated errors from the inversion process.
Recently, attention control (cao2023masactrl; hertz2022prompt) has shown promise in achieving high fidelity and essential content preservation. For instance, MusicMagus (zhang2024musicmagus) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks. Note that in this paper, rigid music editing refers to structural modifications requiring recomposition of core musical relationships, such as the instruments changes, genre transformation, macro-harmonic restructuring, etc; whereas, non-rigid music editing involves parametric adjustments preserving original musical relationships, including adjusting the beat, melody, pitch, rhythm, and other subtle aspects, which generally involves more microscopic editing. Nevertheless, attention control methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 4.
@@
In this work, to bridge this gap, we introduce a zero-shot music editing system MEDIC based on an innovative Disentangled Inversion Control technique, which comprises two components of Harmonized Attention Control and Disentangled Inversion. Cross-attention control (hertz2022prompt) and mutual self-attention control (cao2023masactrl) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively. However, simply combining these two approaches sequentially for music editing can result in sub-optimal performance, particularly in the original dual-branch setup, where it struggles with global attention refinement. To address this issue, we propose Harmonized Attention Control, which unifies cross-attention control and mutual self-attention control by introducing an intermediate branch named Harmonic Branch, designed to progressively modify both rigid and non-rigid attributes in music. Furthermore, we disentangle the diffusion process into triple branches and correct the deviation path caused by CFG in the source branch, which affects the essential content preservation. The other branches remain unchanged to ensure high edit fidelity.
@@
Due to the lack of standardized benchmarks in music editing, we also introduce a new benchmark ZoME-Bench, consisting of 1,100 audio clips in 10 rigorously curated editing categories across rigid and non-rigid tasks. Each entry is carefully assembled, comprising a source prompt, a source audio, a target text prompt, human instruction, and blended words intended for editing. We then extensively evaluate MEDIC and baselines on ZoME-Bench and other datasets. Our contributions can be summarized as follows.
  • We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks. We develop a zero-shot music editing system MEDIC based on DIC.
@@
  • Disentangled Inversion Control includes two critical algorithmic designs. (a) A Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing. (b) Disentangled Inversion Technique is proposed to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information.
  • We introduce ZoME-Bench, a new benchmark for music editing complete with comprehensive evaluation metrics. It consists of 1,100 audio clips categorized into 10 rigorously curated editing tasks, encompassing both rigid and non-rigid tasks.
  • Experimental results on ZoME-Bench indicate that MEDIC outperforms competitive baselines, achieving significant improvements in edit fidelity and essential content preservation, as depicted in Figure 3. Moreover, MEDIC achieves state-of-the-art performance under the variable-length music editing settings of the commonly used MusicDelta dataset.
2. Related Work
Text-based Audio editing. Some prior text-based audio editing studies utilize diffusion models to manipulate audio content according to the target prompt provided (paissan2023audio; plitsis2024investigating; han2023instructme). The two primary challenges (i.e. editing fidelity and essential content preservation) command intense focus.
@@ 2. Related Work
Existing methodologies (novack2024ditto; wu2024music) for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (copet2023simple; agostinelli2023musiclm; chen2024musicldm) that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (paissan2023audio; plitsis2024investigating). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text embeddings for signal reconstruction. Some methods choose to employ both strategies (kawar2023imagic), which further increases the computational load. The third path involves inversion techniques, which typically use DDPM (hubermanspiegelglas2024edit; wu2023latent) or DDIM (song2020denoising; zhang2024musicmagus) inversion strategies to extract diffusion noise vectors that match the source signal. Considering its rapid and intuitive zero-shot editing capabilities, in this work, we choose inversion techniques as our primary research framework. Different from existing inversion strategies, we propose a new inversion technique named Disentangled Inversion Control, which disentangles the diffusion process into triple branches with both mutual self-attention control and cross-attention control to achieve accurate editing while preserving structural information.
@@
Inversion Techniques. The field of image inversion techniques has experienced significant progress in recent years (brooks2023instructpix2pix; kim2022diffusionclip; parmar2023zero; dhariwal2021diffusion). Although DDIM inversion proves to be effective for unconditional diffusion models (song2020denoising; yang2025emovoice; liu2025speech), its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. Various solutions have been proposed to address these challenges (mokady2023null; tumanyan2022plugandplay). For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during editing. In contrast, Edit-Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (hubermanspiegelglas2024edit). Optimization-based inversion methods using specific latent variables have recently gained popularity (ju2023direct; kawar2023imagic). These methods are designed to minimize the accumulated errors that result from the inversion of DDIM. Techniques such as Null-Text Inversion (mokady2023null) are promising, but introduce complexity and instability into the optimization process. Different from these inversion techniques, we introduce a plug-and-play method called Disentangled Inversion Control to separate branches which enables each branch to unleash its maximum potential individually, achieving superior performance with considerably fewer computational resources.
@@
Figure 4. Comparisons between our method and major baselines based on two-branch inversion techniques, including DDIM Inversion (song2020denoising) and MusicMagus (zhang2024musicmagus). (a) Framework of DDIM Inversion, showing configurations with and without classifier-free guidance. (b) Framework of MusicMagus, which incorporates cross-attention control. (c) Framework of our method, featuring Disentangled Inversion Control. (d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch method.
3. Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis. We further analyze the challenges associated with these methods.
3.1. Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition $\bm{c}$} to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following this objective:
[…]
where noise is added to the sampled data ${\bm{z}}_{0}$ according to diffusion time step $T$. During inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained diffusion model for $T$ steps. To generate audio from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling, where $\alpha$ is hyperparameter:
(2) ${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$
3.2. DDIM Inversion
While diffusion models have superior characteristics in the feature space that can support various downstream tasks, it is hard to apply them to audio in the absence of natural diffusion feature space for non-generated audio. Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
(3) ${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$
where $\bm{z}$$^{*}$ denotes DDIM inversion latent. However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$ in Equation 2, Equation 3 and Figure 4(a). Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ as shown in Figure 4(a):
(4) ${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$
3.3. Classifier-free Guidance
Classifier-free Guidance (CFG) (classifier-free) is proposed to overcome the limitation of weak text adherence in text-conditioned models. The modified noise estimation in CFG can be expressed as:
(5) $\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$
where $\varnothing$ is the embedding of a null text. A higher guidance scale $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error. CFG further leads to another perturbation from ${\bm{z}}^{src^{\prime}}_{t}$ to ${\bm{z}}^{src^{\prime\prime}}_{t}$ due to the destruction of the DDIM process and error augmentation, as depicted in Figure 4(a). This issue becomes problematic in editing scenarios where precise control over audio synthesis is desired.
4. Methodology
4.1. Task Definition
@@
Despite significant work in text-to-audio generation models (huang2023makeanaudio; liu2024audiolcm; liu2023audioldm2; liu2024flashaudio), particularly with the emergence of Latent Diffusion Models (LDM) (liu2023audioldm; chen2024musicldm), research on zero-shot music editing remains limited. Given the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt $\mathcal{P}$}, zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to directly modify ${\bm{x}}^{src}_{0}$ and $\mathcal{P}$}, and synthesize the desired music ${\bm{x}}^{tgt}_{0}$, which is aligned with the target edited text prompt $\mathcal{P}$$^{*}$. We compress source audio signal ${\bm{x}}^{src}_{0}$ into latent ${\bm{z}}^{src}_{0}$ for inversion.
4.2. Disentangled Inversion Control